OpenStudy (anonymous):

Which of the following is a solution for the system y = -3x + 5 and y = 2x + 5?

jimthompson5910 (jim_thompson5910):

y = 2x + 5 -3x + 5 = 2x + 5 now solve for x

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (anonymous):

(1, 7) (5, 0) (2, -1) (0, 5) i got c?

jimthompson5910 (jim_thompson5910):

what do you get when you solve -3x + 5 = 2x + 5 for x

jimthompson5910 (jim_thompson5910):

were you able to solve that equation?

jimthompson5910 (jim_thompson5910):

here's a hint -3x + 5 = 2x + 5 -3x - 2x = 5 - 5 -5x = 0